Find the value and interest of $80 deposited at the end of each month at 4.5% compounded monthly for 25 years
Answer by stanbon(75887)

Find the value and interest of $80 deposited at the end of each month at 4.5% compounded monthly for 25 years
-----
Formula:
MP =P[(i/C)(1+(1/C))^(NC))/((1+i/C)^(NC) - 1]
Solve for "P"
-------------------
Note::
i/C 0.045/12 = 0.00375
NC = 25*12 = 300
MP= 80
--------------
80 = P[(0.00375)(1.00375)^300]/(1.00375)^300-1)]
-----
P = 80/[(0.00375)(1.00375)^300]/(1.00375)^300-1)]
---
P = $14392.83 (amt. of the original debt to be paid off)
----------------------
Interest = (total paid - owed)/(owed) = ([(80*12*25)-14392.83]/14392.83
= 0.67 = 67%
-----
Cheers,
Stan H.
---------